OpenStudy (anonymous):

Two boxes of apples contain a total of 24 apples. If you halve the number of apples in the first box and add 4 apples to the second box, the total changes to 20 apples. How many apples are in each box initially? 12 and 12 18 and 6 16 and 8 10 and 14

OpenStudy (anonymous):

An equation in the system is y = x + 1. What is the other equation in the system if the solution is only at (0, 1)? x − y = 1 2x + 2y = 4 x + y = 1 x + 2y = 2

OpenStudy (jadeishere):

so half of 5 + 18 doesn't equal 20 from what i know

OpenStudy (jadeishere):

No, it's not D..

OpenStudy (anonymous):

it don't matter I chose D

OpenStudy (anonymous):

An equation in the system is y = x + 1. What is the other equation in the system if the solution is only at (0, 1)? x − y = 1 2x + 2y = 4 x + y = 1 x + 2y = 2

OpenStudy (jadeishere):

Work out all of the problems

OpenStudy (anonymous):

idk

OpenStudy (jadeishere):

Hold on, let me work it out @catking

OpenStudy (anonymous):

ok thx

OpenStudy (jadeishere):

it's 16 + 8 for the first question

OpenStudy (jadeishere):

the reason being is if you half 16 you get 8 if you add four to 8 you get 12 12 + 8 = 20

OpenStudy (anonymous):

so whats the answer?

OpenStudy (jadeishere):

Okay for the equation... the other equation is perpendicular to y = x + 1 with the intersecting intercepts of (0, 1)?

OpenStudy (jadeishere):

I might need help with this one!! @beccaboo333 @mathslover @arabpride